OpenStudy (anonymous):

i only have 36 days to complete odysseytwere and will fan and medal people who can take my pass code and stuff and do it

OpenStudy (anonymous):

I'm sorry, but I don't think anyone will do your whole test for you

OpenStudy (anonymous):

well help me on some stuff

OpenStudy (anonymous):

Which of the following equations is not equivalent to the others? y = -3x - 4 y = -3x + (-4) y = -4 + 3x y = -4 - 3x